### v2

# Fast Crystal Tensor Property Prediction: A General O(3)-Equivariant Framework Based on Polar Decomposition

Abstract
Predicting the tensor properties of crystalline materials is a fundamental task in materials science. Unlike single-value property prediction, which is inherently invariant, tensor property prediction requires maintaining $O(3)$ group tensor equivariance. This equivariance constraint often introduces tremendous computational costs, necessitating specialized designs for effective and efficient predictions. To address this limitation, we propose a general $O(3)$-equivariant framework for fast crystal tensor property prediction, called GoeCTP. Our framework is efficient as it does not need to impose equivalence constraints onto the network architecture. Instead, GoeCTP captures the tensor equivariance with a simple external rotation and reflection (R&R) module based on polar decomposition. The crafted external R&R module can rotate and reflect the crystal into an invariant standardized crystal position in space without introducing extra computational cost. We show that GoeCTP is general as it is a plug-and-play module that can be smoothly integrated with any existing single-value property prediction framework for predicting tensor properties. Experimental results indicate that GoeCTP achieves higher prediction performance and runs 13$\times$ faster compared to existing state-of-the-art methods in elastic benchmarking datasets, underscoring its effectiveness and efficiency.
1 Introduction
The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering (Yan et al., 2024b). Compared to single-value property prediction, predicting tensor properties of crystalline materials is substantially more complex. This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress (Nye, 1985; Resta, 1994; Yan et al., 2024b). Consequently, tensor property prediction modeling necessitates preserving consistency with the crystal’s spatial position, exhibiting “special” $O(3)$ equivariance (Yan et al., 2024b). Furthermore, high-order tensor property prediction is computationally intensive due to its high dimensionality and large data volume. Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
Thus far, several works have been dedicated to crystal tensor property prediction. One prominent category of approaches involves ab initio physical simulation techniques, such as density functional theory (DFT) (Petousis et al., 2016). These classical simulation techniques can accurately calculate various material properties, including electronic structures, phonon spectra, and tensor properties. However, they necessitate extensive computational resources due to the complexity of handling a vast number of atoms and electrons in crystal systems (Yan et al., 2024b), hindering their applicability in practice. Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction. These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) (Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Xie & Grossman, 2018) or transformers (Yan et al., 2024a; Taniai et al., 2024; Yan et al., 2022; Lee et al., 2024; Wang et al., 2024a).
Existing ML methods have primarily focused on predicting single-value properties of crystals. They exhibit promising prediction performance by enforcing the periodic invariance constraint into the model design. However, these methods cannot be directly applied to tensor property prediction as they overlook the inherent anisotropy of the crystal systems, failing to achieve “special” $O(3)$ equivariance (Yan et al., 2024a; Taniai et al., 2024; Yan et al., 2022; Wang et al., 2024a). Some recent works attempt to ensure equivariance through specialized designs of network architectures (Mao et al., 2024; Lou & Ganose, 2024; Heilman et al., 2024; Wen et al., 2024; Pakornchote et al., 2023; Yan et al., 2024b; Zhong et al., 2023). These methods generally employ harmonic decomposition to achieve equivariance, where the tensor space is decomposed into the direct sum of irreducible representations of the rotation group. In this regard, numerous computationally intensive operations, such as tensor products and the merging of irreducible representations, are required. These operations incur large amounts of computation overhead, particularly when processing high-order data.
To address the above challenges, we propose a novel $O(3)$-equivariant framework for fast crystal tensor property prediction dubbed GoeCTP. In particular, instead of enforcing the equivariance into the model design, we craft a simple yet effective “rotation and reflection” (R&R) module based on polar decomposition. Our designed R&R module rotates and reflects the input crystal with different positions into a standardized invariant position in space. This standardized crystal position is subsequently passed into the property prediction network to obtain an invariant tensor value. Meanwhile, the orthogonal matrix obtained from the R&R module is used to achieve equivariant tensor properties prediction. Our approach is plug-and-play as it can be readily integrated with any existing single-value property prediction network for predicting tensor properties without incurring additional computational costs. Compared to the previous state-of-the-art work Yan et al. (2024b), the GoeCTP method achieves higher quality prediction results and runs more than 13$\times$ faster in the elastic benchmarking dataset.
2 Preliminaries and problem statement
2.1 Preliminaries
The structure of crystalline materials consists of a periodic arrangement of atoms in 3D space, with a repeating unit called a unit cell. An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters (Yan et al., 2022; Wang et al., 2024b; Jiao et al., 2024). General methods for describing crystals can be divided into the Cartesian coordinate system and the fractional coordinate system, as described below.
Cartesian Coordinate System. A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{T}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell. Each $\bm{a}_{i}\in\mathbb{R}^{d_{a}}$ is a $d_{a}$-dimensional feature vector characterizing an individual atom. The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{T}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell. The lattice maxtrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ consists of the lattice vectors $\bm{l}_{1}$, $\bm{l}_{2}$, and $\bm{l}_{3}$, which form the basis of the 3D Euclidean space. A complete crystal is therefore represented as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\bm{a}_{i}},\hat{\bm{x}}_{i})|\hat{\bm{x}}_{i}=\bm{x}_{i}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3},\hat{\bm{a}_{i}}=\bm{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$. In this representation, the integers $k_{i}$ and $l_{i}$ denote all possible atomic positions in the periodic lattice.
Fractional Coordinate System. Instead of using the standard orthogonal basis, fractional coordinate system utilizes the lattice matrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ as the basis vectors for atomic positions. With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{T}\in[0,1)^{3}$. The corresponding Cartesian coordinate vector can then be expressed as $\bm{x}_{i}=f_{i}\bm{l}_{i}$. Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{T}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. Some of the visualizations in the figure are generated using the VESTA (Momma & Izumi, 2011), Yan et al. (2024b), and VELAS (Ran et al., 2023).
2.2 Problem statement
Crystal Tensor Prediction. The crystal tensor properties prediction is a classic regression task. Its goal is to estimate the high-order tensor property denoted as $\mathcal{Y}_{label}$ from the raw crystal data represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$ while minimizing the expected error between the actual property $\mathcal{Y}_{label}$ and the estimated property value $\mathcal{Y}_{pred}$. This problem can be formally formulated as follows:
$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}_{label}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}^{(n)}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset. In what follows, we will omit superscript $n$ for simplicity. As did in the literature (Mao et al., 2024; Wen et al., 2024; Yan et al., 2024b), our objective is to estimate high-order tensor properties, including dielectric tensor (i.e., $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$) and elastic tensor (i.e., $\mathcal{Y}_{label}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
$\mathbf{O(3)}$-Equivariance. The $O(3)$ group, also known as the orthogonal group in 3D, consists of all rotations and reflections in 3D space, i.e., $3\times 3$ orthogonal matrices. In predicting the crystal tensor properties, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies (Hoogeboom et al., 2022; Song et al., 2024). Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq. 1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
$f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{T},$ (2)
where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix (Yan et al., 2024b). For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig. 1.
Our Objective. Our objective is to devise a new framework that can capture the $O(3)$-equivariance for accurately predicting the high-order crystal tensor properties.
Figure 2: The Illustration of GoeCTP. To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary direction, to the standardized crystal position. Next, (2) Crystal Graph Construction module organizes the adjusted input into a crystal graph, followed by (3) the Node & Edge Feature Embedding module, which encodes the features of the crystal graph. Subsequently, (4) the Prediction module leverages these embedded features to predict the invariant tensor properties corresponding to the standardized crystal position. Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the output of the equivariant tensor properties.
3 Methodology
3.1 The Rationale of Our Framework
To enforce $O(3)$ equivariance, as described in Eq. 2, existing crystal tensor prediction methods typically employ specialized network architectures based on harmonic decomposition, resulting in substantial computational overhead. For example, existing work (Yan et al., 2024b; a) gather the rotational information from neighboring nodes to the central node $i$ through the following computation formulated:
$\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})),$ (3)
where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ (Geiger & Smidt, 2022), and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics. When transmitting rotational information for higher-order tensors like the elastic tensor, the rotation order $\lambda$ must be set to 4 (Yan et al., 2024b). In comparison, for transmitting lower-order tensor rotational information (where $\lambda$ is set to 1 (Yan et al., 2024a)), the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})$ increases fourfold, thus, the corresponding computational cost increases. If we can transmit the rotational information for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, this would significantly reduce the computational costs.
To this end, we propose the GoeCTP method, which achieves $O(3)$ equivariance without requiring specialized network architecture design. For instance, it enables the transmission of rotational information for higher-order tensors using $\lambda=1$ rather than $\lambda=4$, thereby simplifying the model complexity while maintaining equivariance. GoeCTP leverages polar decomposition, a mathematical technique with significant geometric properties. Polar decomposition decomposes a matrix into two components: an orthogonal matrix, representing a rotation or reflection, and a positive semi-definite matrix (Higham, 1986). The orthogonal matrix captures an $O(3)$-group transformation, while the positive semi-definite matrix represents a stretching or scaling. When applied to crystal structures, polar decomposition effectively separates the spatial directional information of the crystal via the orthogonal matrix, allowing us to directly transfer equivariance without the need for intricate model design. Furthermore, polar decomposition is only performed during the data preprocessing stage and does not occur during network training, thereby introducing virtually no additional computational burden.
3.2 Our Proposed Framework: GoeCTP
In what follows, we will first introduce the core rotation and reflection (R&R) module of the proposed GoeCTP, focusing on how to obtain a standardized crystal position for a crystal with arbitrary spatial direction using polar decomposition. Then, we will describe how the input crystal data is processed and introduce the property prediction network of GoeCTP. Finally, we will explain how proposed GoeCTP achieves equivariant property predictions. An overview of the GoeCTP framework is illustrated in Fig. 2.
R&R Module. The primary function of the R&R Module is to rotate and reflect the crystal to a standardized crystal position in space. In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for the R&R module to perform rotation and reflection operations on the crystal data. Therefore, R&R module adopts the fractional coordinate system to represent the crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. This choice simplifies the implementation of the R&R module. Specifically, the application of $O(3)$-group transformations is achieved by applying polar decomposition as follows.
Proposition 3.1
(Polar Decomposition (Hall & Hall, 2013; Higham, 1986; Jiao et al., 2024).) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\exp(\mathbf{S})$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{S}\in\mathbb{R}^{3\times 3}$ is a symmetric matrix and $\exp(\mathbf{S})=\sum_{n=0}^{\infty}\frac{\mathbf{S}^{n}}{n!}$ defines the exponential mapping of $\mathbf{S}$.
Building on the above proposition, it is evident that for a crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the invertible lattice matrix $\mathbf{L}$ can be uniquely represented by a symmetric matrix $\mathbf{S}$. By defining $\mathbf{H}=\exp(\mathbf{S})$, the lattice matrix $\mathbf{L}$ can also be uniquely represented by a symmetric matrix denoted as $\mathbf{H}$, i.e. $\mathbf{L}=\mathbf{Q}\mathbf{H}$. Furthermore, $\mathbf{H}$ remains invariant with the $O(3)$ group transformation.
Proof 3.1.1
Given the lattice matrices $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$, where $\mathbf{Q^{\prime}}$ is an arbitrary $O(3)$ group transformation, applying polar decomposition on both $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$ yields the following expressions:
$\mathbf{L}=\mathbf{Q}\mathbf{H}$ (4)
$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (5)
By combining Eq. 4 and 5, we can have the following equation:
$\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (6)
Because $\mathbf{Q^{\prime}}\mathbf{Q}(\mathbf{Q^{\prime}}\mathbf{Q})^{T}=\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{Q}^{T}\mathbf{Q^{\prime}}^{T}=\mathbf{I}$, it follows that $\mathbf{Q^{\prime}}\mathbf{Q}$ is an orthogonal matrix. According to the uniqueness of polar decomposition, we can have $\mathbf{Q^{\prime}}\mathbf{Q}=\mathbf{Q_{1}}$, which implies $\mathbf{Q_{1}}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}$. If $\mathbf{H_{1}}\neq\mathbf{H}$, this would violate the uniqueness property of polar decomposition. Therefore, we conclude that $\mathbf{H_{1}}=\mathbf{H}$, indicating that $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation.
In essence, regardless of the $O(3)$ group transformation applied to the crystal in space, the corresponding $\mathbf{H}$ associated with $\mathbf{L}$ remains invariant. Consequently, $\mathbf{H}$ serves as a fixed standardized crystal position for the crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, allowing the crystal with different position to be consistently rotated to a same position through polar decomposition. The concept of the standardized crystal position can be formally stated as the following proposition :
Proposition 3.2
(Standardized Crystal Position.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$, the standardized crystal position $\mathbf{H}$ is obtained. When an arbitrary $O(3)$ group transformation $\mathbf{Q^{\prime}}$ is applied to a crystal structure, $\mathbf{H}$ remains unchanged.
Therefore, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ formulated as:
$f_{p}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H}),$ (7)
where $f_{p}(\mathbf{M})$ represents the application of polar decomposition to the lattice matrix $\mathbf{L}$ of $\mathbf{M}$.
As shown in Fig. 2, the crystal $(\mathbf{A},\mathbf{F},\mathbf{H})$ is passed to the subsequent crystal graph construction module for further processing. The orthogonal matrix $\mathbf{Q}$ obtained during this decomposition is passed to the reverse R&R module, ensuring the equivariant transformation of the output tensor properties. Our proposed R&R module based on polar decomposition allows the input crystal data to be transformed into a standardized spatial position that is invariant under $O(3)$ space group transformations. With this particular module, the equivariance can be captured, meaning that the subsequent components of GeoCTP are no longer required to account for equivariance.
Crystal Graph Construction. To enable networks to handle such infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is typically necessary to employ graph construction methods that represent the interactions between infinite crystal structures and atoms using finite graph data. Here, we use the crystal graph construction from Yan et al. (2024a; b) to describe the structure and relationships within crystals. Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{P},\mathbf{H})$ as introduced in Sec. 2. We assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$. $\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$. When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e. $d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$. Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$. Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al. (2024a); Yan et al. (2024a).
Node and Edge Feature Embedding. Building on previous work (Xie & Grossman, 2018; Yan et al., 2024b; a), node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$. Edge features $\mathbf{e}_{ij}$ are decomposed into their magnitude $||\mathbf{e}_{ij}||_{2}$ and a normalized direction vector $\hat{\mathbf{e}}_{ij}$. The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding (Lin et al., 2023). Subsequently, $\mathbf{e}_{ij}$ are embedded into feature vector $\bm{f}_{ij}^{e}$.
Prediction Module. Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by Yan et al. (2024a); Taniai et al. (2024); Yan et al. (2022); Lee et al. (2024); Wang et al. (2024a), among others. For better performance, we select eComformer (Yan et al., 2024a), which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module. A detailed explanation of eComFormer can be found in Appendix A.1. Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows:
$\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}.$ (8)
Reverse R&R Module. The primary function of the Reverse R&R module is to generate the equivariant tensor property predictions based on the crystal global features from the Prediction module. First, the Reverse R&R module transforms $\bm{G}^{\mathrm{final}}$ into a tensor output, as follows:
$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}}),$ (9)
$\bm{\varepsilon}^{\mathrm{final}}=f_{rp}(\bm{\varepsilon},\mathbf{Q}),$ (10)
where $f_{MLP}(\cdot)$ represents a multilayer perceptron (MLP) and the operation reshaping dimension. Next, $f_{rp}(\cdot)$ utilizes the orthogonal matrix $\mathbf{Q}$ obtained from the R&R Module to convert the tensor output $\bm{\varepsilon}$ into its final equivariant form denoted as $\bm{\varepsilon}^{\mathrm{final}}$.
This conversion $f_{rp}(\cdot)$ for predicting the dielectric tensor can be expressed by:
$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{T}.$ (11)
For predicting the higher-order elastic tensor, the conversion process becomes more complex. Specifically, the conversion is expressed as follows:
$C_{ijkl}^{\mathrm{final}}=\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{p=1}^{3}\sum_{q=1}^{3}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq},$ (12)
where the subscripts denote the specific positions of the elements within the tensor data.
4 Related work
GNN-Based Methods. CGCNN is a pioneering GNN model specifically designed for handling crystal structures (Xie & Grossman, 2018). This model proposed to represent crystal structures as multi-edge crystal graphs. It was applied to predict various single-value properties such as formation energy and band gap. Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge (Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Das et al., 2023; Lin et al., 2023). These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors. Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction. In contrast, recent studies attempted to ensure equivariance through specialized network architectures (Mao et al., 2024; Lou & Ganose, 2024; Heilman et al., 2024; Wen et al., 2024; Pakornchote et al., 2023; Yan et al., 2024b; Zhong et al., 2023). These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties. Within these network architectures, many operations are required, such as tensor products and combining irreducible representations. These processes significantly increase computational costs, especially when handling higher-order data.
Transformer-Based Methods. Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties. Matformer (Yan et al., 2022), one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells. This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures. Subsequently, more advanced Transformer-based approaches were proposed. Some methods, such as ComFormer (Yan et al., 2024a), CrystalFormer (Wang et al., 2024a), and CrystalFormer (Wang et al., 2024a), typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties. Furthermore, DOSTransformer (Lee et al., 2024) is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance. Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
5 Experiments
5.1 Experimental setup
Datasets. In this work, we evaluate the performance of GoeCTP on two key tensor property prediction tasks: the second-order dielectric tensor and the fourth-order elastic tensor, respectively. The dataset for dielectric tensor is derived from the data processed by Yan et al. (2024b), sourced from the JARVIS-DFT database. Similarly, the dataset for elastic tensor is obtained from the dft_3d data within the JARVIS-DFT database111 https://pages.nist.gov/jarvis/databases/. The statistical details of the datasets are presented in Table 1. Additionally, in the dielectric tensor dataset, the dielectric tensor is a $3\times 3$ symmetric matrix. Therefore, during prediction, we predict 6 elements of the matrix (see Appendix A.2 for details) and then reconstruct the entire $3\times 3$ symmetric matrix. In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor (Itin & Hehl, 2013; Wen et al., 2024). Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.2 for details).
Dataset Sample size Fnorm Mean Fnorm STD Unit
Dielectric 4713 14.7 18.2 Unitless
Elastic 25110 306.4 238.4 GPa
Table 1: Dataset statistics.
Baseline Methods. we selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e. MEGNET (Chen et al., 2019; Morita et al., 2020), EGTNN (Zhong et al., 2023), and GMTNet (Yan et al., 2024b), as baseline methods.
[…]
Experimental Settings. The experiments were performed using an NVIDIA GeForce RTX 3090 GPU. For benchmarking, we directly utilized the codebases for MEGNET, ETGNN, and GMTNet as provided by Yan et al. (2024b). For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio. During model training, Huber loss, AdamW (Loshchilov & Hutter, 2018), a weight decay of $10^{-5}$, and polynomial learning rate decay were employed.
Since we observed some randomness in the results during the training process of the repeated dielectric dataset experiments, we conducted 5 repeated experiments for each method and used the average of the 5 experiment metrics as the final experimental results. The results of the 5 repeated experiments can be found in the Appendix 11.
5.2 Experimental Results
Predicting Dielectric Tensors. The performance of various models in predicting the dielectric tensor is summarized in Table 2. While GoeCTP and GMTNet show close results in terms of Fnorm and EwT 25%, GoeCTP demonstrates higher values for EwT 10% and EwT 5%, with an improvement of 5% in EwT 5% compared to GMTNet. This indicates that GoeCTP delivers higher-quality predictions compared to GMTNet. Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer (Yan et al., 2024a) and CrystalFormer (Taniai et al., 2024). Detailed results of these combinations are provided in Appendix 11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5%
Table 2: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the dielectric dataset.
Predicting Elastic Tensors. Experimental results on the elastic tensor dataset are presented in Table 3. GoeCTP demonstrated outstanding performance in predicting higher-order, complex tensors, outperforming all baseline methods across every evaluation metric. Notably, it achieved the lowest Fnorm of 107.11 and improved all EwT metrics by an average of 6% compared to GMTNet, emphasizing its general applicability in predicting diverse tensor properties in materials science. Further details on the results of GoeCTP’s combination with other models can be found in Appendix 11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2%
Table 3: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the elastic dataset.
Verifying the $O(3)$ Equivariance. To evaluate the effectiveness of GoeCTP, we conducted experiments to verify the $O(3)$ equivariance of tensor properties. Specifically, after training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set). All crystals in the original test set were adjusted to the standardized crystal position, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set. The method for generating the corresponding orthogonal matrices is from Heiberger (1978). We then evaluated both the Prediction module (eComFormer) and GoeCTP on these two datasets and compared the performance metrics.
The results on the dielectric tensor dataset are shown in Table 4. GoeCTP performed equally well on the augmented test set as on the original test set, indicating that it maintains strong $O(3)$ equivariance for tensor properties. In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45%, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties. The results on the elastic tensor dataset are shown in Table 5. Similar to the dielectric tensor dataset, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm dropping by nearly 29%. In contrast, GoeCTP’s performance remained unchanged, further demonstrating the effectiveness of the GoeCTP method. Additionally, runtime comparisons between GoeCTP and eComFormer on the test set showed that GoeCTP introduced no significant increase in runtime. This indicates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality without compromising efficiency.
eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
Fnorm $\downarrow$ 3.23 4.71 3.23 3.23
EwT 25% $\uparrow$ 83.2% 69.7% 83.2% 83.2%
EwT 10% $\uparrow$ 56.8% 42.7% 56.8% 56.8%
EwT 5% $\uparrow$ 35.5% 22.5% 35.5% 35.5%
Total Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric dataset.
eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
Fnorm $\downarrow$ 107.11 138.45 107.11 107.11
EwT 25% $\uparrow$ 42.5% 25.9% 42.5% 42.5%
EwT 10% $\uparrow$ 15.3% 2.2% 15.3% 15.3%
EwT 5% $\uparrow$ 7.2% 0.2% 7.2% 7.2%
Total Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
Table 5: Ablation study for verifying the $O(3)$ equivariance with elastic dataset.
Efficiency. The results presented in Table 6 and Table 7 illustrate the running times for GoeCTP and baseline methods. On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet. On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process. To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply. In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a multilayer perceptron (MLP) at the network’s output, ensuring both efficiency and scalability across different tensor orders.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 663 1325 1611 616
Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048
Table 6: Efficiency comparison on the dielectric dataset.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422
Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189
Table 7: Efficiency comparison on the elastic dataset.
6 Conclusion, limitations, and future works
In this work, we propose a novel $O(3)$-equivariant framework GoeCTP for fast crystal tensor prediction, which is a plug-and-play framework that can be integrated with any existing single-value property prediction network, enabling them to predict tensor properties with almost no additional computational cost. Based on predicting invariant tensor properties from standardized crystal positions and using an external module to ensure tensor equivariance, this approach has achieved state-of-the-art performance and the highest efficiency in different widely used crystal tensor property benchmark tests. The limitations of our current GoeCTP include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module and, in theory, does not introduce additional performance enhancements. (2) Currently, GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types. In future work, we intend to explore the following directions: for (1), we aim to improve performance by integrating prior knowledge related to the independent components of tensor properties across different crystal systems in Appendix A.2; for (2), we plan to extend the framework to other domains, such as 3D point cloud research, broadening its applicability.
Appendix A Appendix
A.1 Details of eComFormer
The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks (Yan et al., 2024a). It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph. The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. (2024a).
[…]
$\bm{\alpha}_{ij}=\frac{\bm{q}_{ij}\circ\bm{\xi}_{K}(\bm{k}_{ij})}{\sqrt{d_{\bm{q}_{ij}}}},\bm{msg}_{ij}=\operatorname{sigmoid}(\operatorname{BN}(\bm{\alpha}_{ij}))\circ\bm{\xi}_{V}(\bm{\upsilon}_{ij}),$ (13)
where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively, and the operators $\circ$ denote the Hadamard product, $\bm{\xi}_{msg}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$. Then, node feature $\bm{f}_{i}$ is updated as follows,
$\bm{m}s\bm{g}_{i}=\sum_{j\in\mathcal{N}_{i}}\bm{m}s\bm{g}_{ij},\bm{f}_{i}^{\mathrm{new}}=\bm{\xi}_{msg}(\bm{f}_{i}+\mathrm{BN}(\bm{m}s\bm{g}_{i})),$ (14)
where $\bm{\xi}_{msg}(\cdot)$ denoting the softplus activation function.
The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers (Geiger & Smidt, 2022). It uses node feature $\bm{f}_{i}^{l}$ and equivalent vector feature $\mathbf{e}_{ji}$ embedded by corresponding spherical harmonics $\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})=c_{0}$,$\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})=c_{1}*\frac{\mathbf{e}_{ji}}{||\mathbf{e}_{ji}||_{2}}\in\mathbb{R}^{3}$ and $\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{5}$ to represent the input features. Gathering rotational information from neighboring nodes to the central node $i$, the first TP layer is shown as
$\bm{f}_{i,0}^{l}=\bm{f}_{i}^{l^{\prime}}+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{0}(\hat{\mathbf{e_{ji}}})),\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})),\lambda\in\{1,2\},$ (15)
where $\bm{f}_{i}^{l^{\prime}}$ is linearly transformed from $\bm{f}_{i}^{l}$, $|\mathcal{N}_{i}|$ and $\mathbf{TP}_{\lambda}$ represent the number of neighbors of node $i$ and TP layers with rotation order $\lambda$ respectively. Then, to represent the invariant node features, the second TP layer is written as
$\bm{f}_{i}^{l*}=\frac{1}{|\mathcal{N}_{i}|}(\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,0}^{l},\mathbf{Y}_{0}(\hat{\mathbf{e_{ji}}}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,1}^{l},\mathbf{Y}_{1}(\hat{\mathbf{e_{ji}}}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,2}^{l},\mathbf{Y}_{2}(\hat{\mathbf{e_{ji}}}))),$ (16)
stacking the two tensor product layers together using both linear and nonlinear transformations, the output $\bm{f}_{i,updated}^{l}$ is combined as
$\bm{f}_{i,updated}^{l}=\bm{\sigma}_{\mathrm{equi}}(\mathrm{BN}(\bm{f}_{i}^{l*}))+\mathrm{LN}_{\mathrm{equi}}(\bm{f}_{i}^{l}),$ (17)
with $\bm{\sigma}_{\mathrm{equi}}$ denoting a nonlinear transformation made up of two softplus layers with a linear layer positioned between them. The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer are shown in Fig. 3.
A.2 Tensor properties symmetry
Tensor properties, such as dielectric tensors and elastic tensors, describe the material’s response to external physical fields (such as electric fields or mechanical stress). In materials with symmetry, this response must adhere to the symmetry requirements of the material. As demonstrated by Yan et al. (2024b), when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$,the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$. Therefore, the corresponding tensor properties also remain unchanged, i.e., $\bm{\varepsilon}=\mathbf{R}\bm{\varepsilon}\mathbf{R}^{T}$. Thus, crystal symmetry imposes strict constraints on the components of the tensor, leading to the simplification or elimination of many components, reducing the number of independent components.
Crystal system Number of independent elements Dielectric tensor
Cubic 1 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Tetragonal & Hexagonal & Trigonal 2 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Orthorhombic 3 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{22}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Monoclinic 4 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&\bm{\varepsilon}_{13}\\ 0&\bm{\varepsilon}_{22}&0\\ \bm{\varepsilon}_{13}&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Triclinic 6 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&\bm{\varepsilon}_{12}&\bm{\varepsilon}_{13}\\ \bm{\varepsilon}_{12}&\bm{\varepsilon}_{22}&\bm{\varepsilon}_{23}\\ \bm{\varepsilon}_{13}&\bm{\varepsilon}_{23}&\bm{\varepsilon}_{33}\end{pmatrix}$
Table 8: Number of independent components in the dielectric tensor for different crystal systems.
Independent components in the dielectric tensor. The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry. The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 8 (Mao et al., 2024).
Voigt notation for elastic tensor. Similar to the dielectric tensor, the elastic tensor also has independent elements for various types of systems due to the crystal symmetry. The elastic tensor has a minimum of 3 and a maximum of 21 independent elements for various types of systems. Voigt notation is a compact way to represent these independent components of tensor properties (Itin & Hehl, 2013). According to the rules $11\to 1\mathrm{~{};~{}}22\to 2\mathrm{~{};~{}}33\to 3\mathrm{~{};~{}}23,~{}32\to 4\mathrm{~{};~{}}31,~{}13\to 5\mathrm{~{};~{}}12,~{}21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix (Wen et al., 2024; Ran et al., 2023):
$C=\begin{pmatrix}C_{1111}&C_{1122}&C_{1133}&C_{1123}&C_{1131}&C_{1112}\\ C_{1122}&C_{2222}&C_{2233}&C_{2223}&C_{2231}&C_{2212}\\ C_{1133}&C_{2233}&C_{3333}&C_{3323}&C_{3331}&C_{3312}\\ C_{1123}&C_{2223}&C_{3323}&C_{2323}&C_{2331}&C_{2312}\\ C_{1131}&C_{2231}&C_{3331}&C_{2331}&C_{3131}&C_{3112}\\ C_{1112}&C_{2212}&C_{3312}&C_{2312}&C_{3112}&C_{1212}\end{pmatrix}\to\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}$
The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to Wen et al. (2024); Ran et al. (2023)).
Crystal system Number of independent elements Elastic tensor
Cubic 3 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{12}&0&0&0\\ C_{12}&C_{11}&C_{12}&0&0&0\\ C_{12}&C_{12}&C_{11}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{44}&0\\ 0&0&0&0&0&C_{44}\end{pmatrix}\end{gathered}$
Tetragonal 6 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&0&0&0\\ C_{12}&C_{11}&C_{13}&0&0&0\\ C_{13}&C_{13}&C_{33}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{44}&0\\ 0&0&0&0&0&C_{66}\end{pmatrix}\end{gathered}$
Orthorhombic 9 $C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&0&0&0\\ C_{12}&C_{22}&C_{23}&0&0&0\\ C_{13}&C_{23}&C_{33}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{55}&0\\ 0&0&0&0&0&C_{66}\end{pmatrix}$
Triclinic 21 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}\end{gathered}$
Table 9: Number of independent components in the elastic tensor for different crystal systems.
A.3 Experimental details and additional results
Hyperparameter settings of GoeCTP. When constructing the crystal graph, we used the 16th nearest atom to determine the cutoff radius. For edge embeddings, we used an RBF kernel with $c=0.75$ and values ranging from --4 to 0, which maps $-c/||\mathbf{e}_{ij}||_{2}$ to a 512-dimensional vector. In the dielectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector. For the prediction module, eComFormer, in the dielectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer. For both dielectric and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64. In the Reverse R&R module, for the dielectric tensor prediction task, the 128-dimensional vector features output by the prediction module are mapped to a 6-dimensional vector through two non-linear layers, then reconstructed into a $3\times 3$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ obtained from the R&R module to achieve $O(3)$ equivariant output. For the elastic tensor prediction task, the 256-dimensional vector features output by the prediction module are mapped to a 36-dimensional vector through two non-linear layers, then reconstructed into a $6\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
1st 2nd 3rd 4th 5th mean
MEGNET
Fnorm $\downarrow$ 3.62 3.70 3.66 3.74 3.81 3.71
EwT 25% $\uparrow$ 76.0% 77.1% 75.6% 73.7% 76.4% 75.8 %
EwT 10% $\uparrow$ 41.2% 39.3% 38.2% 39.7% 36.3% 38.9%
EwT 5% $\uparrow$ 17.6% 16.8% 18.5% 19.1% 17.8% 18.0%
ETGNN
Fnorm $\downarrow$ 3.42 3.48 3.33 3.35 3.41 3.40
EwT 25% $\uparrow$ 82.1% 81.7% 83.4% 82.6% 83.2% 82.6 %
EwT 10% $\uparrow$ 46.9% 47.8% 50.7% 49.4% 50.5% 49.1%
EwT 5% $\uparrow$ 22.3% 26.5% 27.1% 25.9% 24.6% 25.3%
GMTNet
Fnorm $\downarrow$ 2.98 3.64 3.23 3.16 3.39 3.28
EwT 25% $\uparrow$ 83.4% 81.3% 83.8% 84.5% 83.7% 83.3 %
EwT 10% $\uparrow$ 54.9% 56.2% 55.8% 56.2% 57.1% 56.0%
EwT 5% $\uparrow$ 31.2% 30.5% 28.8% 29.3% 32.7% 30.5%
GoeCTP (Ours)
Fnorm $\downarrow$ 3.33 3.29 3.36 3.25 2.92 3.23
EwT 25% $\uparrow$ 82.5% 83.4% 82.1% 83.9% 84.2% 83.2 %
EwT 10% $\uparrow$ 57.5% 53.9% 56.4% 56.7% 59.7% 56.8%
EwT 5% $\uparrow$ 36.9% 31.6% 37.1% 34.2% 37.8% 35.5%
Table 10: Comparison of 5 repeated experiments on dielectric dataset.
Hyperparameter settings of GMTNet, ETGNN, and MEGNET. Following Yan et al. (2024b), we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks. The same polynomial learning rate decay scheduler is used in all experiments.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23 3.40 3.53
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2% 81.7% 80.1%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8% 53.8% 52.9%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5% 32.3% 30.6%
Total Time (s) $\downarrow$ 663 1325 1611 616 535 645
Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048 0.042 0.202
Table 11: Additional comparison of performance metrics on dielectric dataset.
Additional results. Given the relatively small size of the dielectric tensor dataset, along with other influencing factors, we observed some randomness in the results during repeated training runs, introducing noise into the outcome comparisons. To mitigate this effect and ensure the reliability of the results, we conducted five independent trials for each method, using the average of the performance metrics across these five trials as the final reported value. The detailed results from these repeated experiments are presented in Table 10.
We assessed the performance of GoeCTP when combined with iComFormer (Yan et al., 2024a) and CrystalFormer (Taniai et al., 2024) (denoted as GoeCTP (iCom.) for the combination) on both the dielectric tensor and elastic tensor datasets. As presented in Table 11, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). In contrast, as shown in Table 12, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11 102.80 107.44
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5% 46.7% 43.5%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3% 18.6% 15.8%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2% 8.2% 7.9%
Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422 4035 7891
Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189 0.315 0.616
Table 12: Additional comparison of performance metrics on elastic dataset.