### v3

# Revisiting the Canonicalization for Fast and Accurate Crystal Tensor Property Prediction

Abstract
Predicting the tensor properties of crystalline materials is a fundamental task in materials science. Unlike single-value property prediction, which is inherently invariant, tensor property prediction requires maintaining $O(3)$ group tensor equivariance. Such equivariance constraint often requires specialized architecture designs to achieve effective predictions, inevitably introducing tremendous computational costs. Canonicalization, a classical technique for geometry, has recently been explored for efficient learning with symmetry. In this work, we revisit the problem of crystal tensor property prediction through the lens of canonicalization. Specifically, we demonstrate how polar decomposition, a simple yet efficient algebraic method, can serve as a form of canonicalization and be leveraged to ensure equivariant tensor property prediction. Building upon this insight, we propose a general $O(3)$-equivariant framework for fast and accurate crystal tensor property prediction, referred to as GoeCTP. By utilizing canonicalization, GoeCTP achieves high efficiency without requiring the explicit incorporation of equivariance constraints into the network architecture. Experimental results indicate that GoeCTP achieves the best prediction accuracy and runs up to 13$\times$ faster compared to existing state-of-the-art methods in benchmarking datasets, underscoring its effectiveness and efficiency.
1 Introduction
The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering [51]. These tensor properties span various orders, such as dielectric tensor with two orders, piezoelectric tensor with three orders, and elastic tensor with four orders. Accurate prediction of these tensor properties is critical for novel materials discovery and design with targeted characteristics. Thus far, several works have been dedicated to crystal tensor property prediction. One prominent category involves ab initio physical simulation techniques, such as density functional theory (DFT) [35]. These classical simulation techniques can provide acceptable error margin for predicting various material properties. However, they necessitate extensive computational resources due to the complexity of handling large crystals with a vast number of atoms and electrons [51], hindering their applicability in practice.
Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction. These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) [1, 26, 2, 47] or transformers [50, 42, 49, 20, 44]. Most existing methods are designed for scalar property prediction, focusing on achieving $SO(3)$ invariance of the crystal structures. However, these scalar-property methods could not be adapted to predict tensor properties of crystalline materials, which exhibits significantly higher complexity. This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress [33, 38, 51]; their modeling necessitates preserving consistency with the crystal’s spatial orientation, exhibiting a unique tensor $O(3)$ equivariance [51].
Therefore, a few recent studies attempt to ensure equivariance through specialized designs of network architectures [28, 25, 11, 46, 34, 51, 53]. These methods generally employ harmonic decomposition to achieve equivariance, where the tensor space is decomposed into the direct sum of irreducible representations of the rotation group. In this regard, numerous computationally intensive operations, such as tensor products and the merging of irreducible representations, are required. These operations incur large amounts of computation overhead, particularly when processing high-order data. Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
One approach that has show promising results in efficient learning with symmetry across various fields, including point clouds [22], n-body simulation [17], and antibodies generation [29], is the use of “canonicalization". Specifically, canonicalization maps a geometric data to an invariant representation [27, 6], referred to as the canonical form, and subsequently reconstructs an equivariant output from the canonical form without imposing any architectural constraints on the network. However, to date, it has not been explored for crystal tensor property prediction.
In this work, we revisit the task of crystal tensor property prediction through the lens of canonicalization and introduce a novel canonicalization strategy tailored for this particular setting. Specifically, rather than embedding equivariance directly into model architecture, we propose a simple yet effective canonicalization module, termed R&R. In particular, our canonicalization is instantiated based on polar decomposition, a continuous mapping technique [6] that can provide enhanced robustness. During prediction, the R&R module applies rotations and reflections to transform the input crystal structure into its canonical form. The canonical form is then fed into a property prediction network to obtain the corresponding canonical tensor representation. Simultaneously, the orthogonal matrix derived from the R&R module is utilized to recover the equivariant output via the tensor transformation rule, enabling equivariant tensor property prediction without incurring additional computational overhead. We conducted experiments on dielectric, piezoelectric, and elastic tensor datasets, respectively, showcasing that the proposed method can achieve $O(3)$-equivariant tensor prediction while maintaining high efficiency. Compared to the previous state-of-the-art work Yan et al. [51], the GoeCTP method achieves higher prediction accuracy and runs by up to 13$\times$ faster in benchmarking datasets.
2 Background
2.1 Preliminaries
The structure of crystalline materials consists of a periodic arrangement of atoms in 3D space, with a repeating unit called a unit cell. An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters [49, 45, 16]. General methods for describing crystals can be divided into the Cartesian coordinate system and the fractional coordinate system, as described below.
Cartesian Coordinate System. A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{\top}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell. Each $\bm{a}_{i}\in\mathbb{R}^{d_{a}}$ is a $d_{a}$-dimensional feature vector characterizing an individual atom. The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell. The lattice maxtrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ consists of the lattice vectors $\bm{l}_{1}$, $\bm{l}_{2}$, and $\bm{l}_{3}$, which form the basis of the 3D Euclidean space. A complete crystal is therefore represented as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\bm{a}_{i}},\hat{\bm{x}}_{i})|\hat{\bm{x}}_{i}=\bm{x}_{i}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3},\hat{\bm{a}_{i}}=\bm{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$. In this representation, the integers $k_{i}$ and $l_{i}$ denote all possible atomic positions in the periodic lattice.
Fractional Coordinate System. Instead of using the standard orthogonal basis, fractional coordinate system utilizes the lattice matrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ as the basis vectors for atomic positions. With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{\top}\in[0,1)^{3}$. The corresponding Cartesian coordinate vector can then be expressed as $\bm{x}_{i}=f_{i}\bm{l}_{i}$. Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. Some of the visualizations in the figure are generated using the VESTA [30], Yan et al. [51], and VELAS [37].
2.2 Problem statement
Crystal Tensor Prediction. The crystal tensor properties prediction is a classic regression task. Its goal is to estimate the high-order tensor property denoted as $\mathcal{Y}$ from the raw crystal data represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. Given that the crystal data $\mathbf{M}$ resides within the input space $\mathcal{V}$, and the tensor property $\mathcal{Y}$ belongs to the separate output space $\mathcal{W}$, the objective of crystal tensor prediction is to find a function ${f_{\theta}}:\mathcal{V}\to\mathcal{W}$ that accurately maps input crystal data to the desired tensor property. This is achieved by minimizing the discrepancy between the true property $\mathcal{Y}$ and the predicted property value $\mathcal{Y}_{pred}$. Therefore, the crystal tensor property prediction task can be mathematically formulated as the following optimization problem:
$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset. In what follows, we will omit superscript $n$ for simplicity. As did in the literature [28, 46, 51], our objective is to estimate high-order tensor properties, including 2-order dielectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$), 3-order piezoelectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$), and 4-order elastic tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
$\mathbf{O(3)}$ group. The $O(3)$ group is an orthogonal group, consisting of rotations and reflections. The elements $g\in O(3)$ act on vectors or tensors in $\mathcal{V}$, $\mathcal{W}$ through their respective group representations, $\rho_{\mathcal{V}}:O(3)\to GL(\mathcal{V})$ and $\rho_{\mathcal{W}}:O(3)\to GL(\mathcal{W})$, where $GL(\mathcal{V})$ and $GL(\mathcal{W})$ are the space of invertible linear maps $\mathcal{V}\to\mathcal{V}$ and $\mathcal{W}\to\mathcal{W}$, respectively. Specifically, the action of an element $g\in O(3)$ on crystal data $\mathbf{M}$ is defined as: $g\cdot\mathbf{M}=\rho_{\mathcal{V}}(g)\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})$. For dielectric tensor property $\mathcal{Y}$, the action of $g$ is given by: $g\cdot\mathcal{Y}=\rho_{\mathcal{W}}(g)\mathcal{Y}=\mathbf{Q}\mathcal{Y}\mathbf{Q}^{\top}$, where $\mathbf{Q}$ is a $3\times 3$ orthogonal matrix (for additional details regarding the transformation of tensor properties, please refer to Appendix A.8).
$\mathbf{O(3)}$-Equivariance for Crystal Tensor Prediction. In the crystal tensor properties prediction task, due to the difference between $\rho_{\mathcal{V}}(g)$ and $\rho_{\mathcal{W}}(g)$, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies [13, 48, 52, 40]. Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq. 1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
$f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top},$ (2)
where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix [51]. For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig. 1 (for more equivariance details, see Appendix A.8).
Figure 2: The Illustration of GoeCTP. To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary orientation, to the canonical form of this crystal. Next, (2) Crystal Graph Construction module organizes the adjusted input into a crystal graph, followed by (3) the Node & Edge Feature Embedding module, which encodes the features of the crystal graph. Subsequently, (4) the Prediction module leverages these embedded features to predict the canonical form of tensor properties corresponding to the canonical form of input crystal. Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the equivariance of the output tensor properties.
3 Methodology
In this section, we will first provide the motivation driving our new framework. We then reformulate the $O(3)$ equivariant tensor prediction task from the perspective of canonicalization. Building upon this foundational perspective, we demonstrate how polar decomposition can be employed as a canonicalization strategy and effectively applied to solve tensor prediction tasks with both effectiveness and efficiency.
3.1 The Motivation of Our Framework
To enforce $O(3)$ equivariance, as described in Eq. 2, existing crystal tensor prediction methods typically employ specialized network architectures based on harmonic decomposition, resulting in substantial computational overhead. For example, existing work gather the rotational information from the neighboring nodes to the central node $i$ through the following computation formulated [51]:
$\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),$ (3)
where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ [7], and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics. To achieve $O(3)$ equivariance of the model and ensure high prediction accuracy for high-order tensor prediction tasks, such as elastic tensor prediction, the rotation order $\lambda$ is set to 3 [51], i.e., $\mathbf{Y}_{3}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{7}$. In contrast, for lower-order tensor prediction tasks, setting $\lambda$ to 1 is sufficient to achieve the model’s $O(3)$ equivariance, i.e., $\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{3}$. Compared to lower-order tensor prediction, the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ nearly doubles for high-order tensor predictions, leading to a significant increase in computational complexity. Consequently, this model architecture introduces tremendous computational costs for tensor property predictions. If we can achieve the model’s $O(3)$ equivariance and maintain the model performance for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, or even without relying on harmonic decomposition, computational costs could be drastically reduced. Inspired by the recent exploration of using canonicalization in efficient learning with symmetry, we aim to propose a principled approach based on canonicalization to enable $O(3)$-equivariant tensor predictions for crystal materials.
3.2 $O(3)$ Equivariant Tensor Prediction from the Perspective of Canonicalization
To mathematically describe $O(3)$-equivariant tensor prediction from the perspective of canonicalization, we first introduce and extend several fundamental concepts derived from canonicalization techniques [27, 17, 6].
Definition 3.1.
(Orbit.) The orbit of a crystal $\mathbf{M}$ is defined as $\text{Orbit}(\mathbf{M})=\{g\cdot\mathbf{M}\mid g\in O(3)\}$.
In the context of crystal tensor prediction, all configurations of a crystal (i.e. a crystal wirh different orientations) in the space fall into the same orbit under the $O(3)$ group transformation. Specifically, for any two crystal configurations denoted as $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ within the same orbit, there exists a rotation or reflection $g\in O(3)$ such that $\mathbf{M}_{1}=g\cdot\mathbf{M}_{2}$. Likewise, the tensor properties of the crystal also exhibit an orbit, i.e. $\text{Orbit}(\mathcal{Y})=\{g\cdot\mathcal{Y}\mid g\in O(3)\}$. All configurations of a crystal within an orbit could be transformed into one another through group transformations. Consequently, we can select a specific configuration from the orbit to serve as a canonical representative (referred to as the canonical form) that effectively characterizes the entire orbit. The formal process is outlined in the following definition.
Definition 3.2.
(Orbit Canonicalization.) A function denoted as $C_{\mathbf{M}}:\mathcal{V}\to\mathcal{V}$ is defined as an orbit canonicalization, if $\forall\mathbf{M}_{1}\in\text{Orbit}(\mathbf{M})$ and $\exists\mathbf{M}_{0}\in\text{Orbit}(\mathbf{M}),$ such that $C_{\mathbf{M}}(\mathbf{M}_{1})=\mathbf{M}_{0}$, where $\mathbf{M}_{0}$ is the canonical form of the orbit.
Orbit canonicalization enables the transformation of a crystal with arbitrary orientations in the space into one with a specific, well-defined orientation. Since the output of orbit canonicalization can always be mapped to the invariant canonical form, it follows that $C_{\mathbf{M}}(\cdot)$ itself is $O(3)$-invariant. Therefore, for any prediction function ${f_{\theta}}(\cdot)$, it holds that ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ remains $O(3)$-invariant (See the proof in Appendix A.1).
Although an $O(3)$-invariant prediction function has been established via orbit canonicalization, the complete process of equivariant prediction requires specific group transformations that map all crystal configurations within the orbit to their canonical form. To address this requirement, we introduce the concept of rigid registration. Rigid registration is a fundamental concept in fields such as computer vision and computational geometry [41]. It aims to find a transformation that matches the position and orientation of one object with the corresponding position and orientation of another object. This concept can be naturally adapted to the crystal, as follows.
Definition 3.3.
(Rigid Registration.) A function $R_{\mathbf{M}}:\mathcal{V}\times\mathcal{V}\to O(3)$ is a rigid registration if $\forall\mathbf{M}_{1},\mathbf{M}_{2}\in\text{Orbit}(\mathbf{M}),R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})=g$, such that $\mathbf{M}_{1}=g\cdot\mathbf{M}_{2}$.
Using Definition 3.3, we can identify a transformation that matches two configurations of a crystal. Specifically, this allows us to find specific group transformations that map all crystal configurations within the orbit to their canonical form. By leveraging Definition 3.2 and Definition 3.3 together, we derive a view for performing $O(3)$-equivariant tensor prediction through canonicalization, as described below.
Proposition 3.4.
($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function formulated as $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$-equivariant for tensor prediction.
The proof of Proposition 3.4 is provided in Appendix A.6. Specifically, in this unified view of $O(3)$ equivariant tensor prediction, we leverage rigid registration, as defined in Definition 3.3, to find specific group transformations that map crystal configurations within the orbit to their canonical form. This canonical form is determined through orbit canonicalization, as described in Definition 3.2. Subsequently, we can directly apply these transformations to the network’s output, thereby effectively achieving the $O(3)$ equivariance. That is, under this perspective, there are no specific requirements or special treatments needed for the network design, as the $O(3)$ transformations applied to the crystal can be externally determined and simultaneously applied directly to the output. Building upon Proposition 3.4, the key challenge in achieving $O(3)$-equivariant tensor prediction lies in identifying the appropriate orbit canonicalization and rigid registration functions.
3.3 Our Proposed Framework: GoeCTP
In what follows, we demonstrate that polar decomposition can effectively serve as both orbit canonicalization and rigid registration functions for crystal tensor prediction. Building on this foundation, we will first introduce the core rotation and reflection (R&R) module of the proposed GoeCTP, which is to obtain the canonical form for a crystal with arbitrary spatial orientations using polar decomposition. Then, we will describe how the input crystal data is processed and introduce the property prediction network of GoeCTP. Finally, we will explain how proposed GoeCTP achieves equivariant tensor predictions. An overview of the GoeCTP framework is illustrated in Fig. 2.
Theorem 3.5.
(Polar Decomposition [9, 12, 16].) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\mathbf{H}$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{H}\in\mathbb{R}^{3\times 3}$ is a Hermitian positive semi-definite matrix.
In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for rotating the crystal configuration to the canonical form. Therefore, throughout this work, we adopts the fractional coordinate system to represent the crystal represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. Then building on Theorem 3.5 , we can define both orbit canonicalization and rigid registration functions for crystal tensor prediction as follows.
Proposition 3.6.
(Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix. In this perspective, $f_{p1}(\cdot)$ serves as the orbit canonicalization function, while $f_{p2}(\cdot)$ serves as the rigid registration function. The crystal configuration $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$ is thus identified as the canonical form.
The detailed proof of Proposition 3.6 is provided in Appendix A.7. Specifically, since the polar decomposition of a lattice matrix always exists and it is unique, any orthogonal matrix acting on the lattice matrix can be separated through polar decomposition, yielding a unique $\mathbf{H}$. Thus, $\mathbf{H}$ can serve as the canonical form. Moreover, the rationale of using polar decomposition is that it is a continuous canonicalization, which can provide improved robustness [6]. For clarity, we provide details on this continuity in Appendix A.14, and further discussions on constructing alternative canonical forms based on polar decomposition in Appendix A.15, respectively. Building upon Proposition 3.6, we utilize polar decomposition to construct the proposed GoeCTP model.
R&R Module. The primary function of the R&R Module is to rotate the crystal configuration to its canonical form and obtain the corresponding transformation. Therefore, based on Proposition 3.6, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ to construct the rigid registration and the orbit canonicalization. Formally, the R&R Module can be expressed as:
$f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H}),\text{ }f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q},$ (4)
where $f_{p1}(\cdot)$ is the orbit canonicalization function, while $f_{p2}(\cdot)$ is the rigid registration function. As shown in Fig. 2, the crystal configuration $(\mathbf{A},\mathbf{F},\mathbf{H})$ derived from orbit canonicalization is passed to the subsequent crystal graph construction module for further processing. The orthogonal matrix $\mathbf{Q}$ obtained during this decomposition is passed to the reverse R&R module, ensuring the equivariant transformation of the output tensor properties. Our proposed R&R module based on polar decomposition allows the input crystal data to be transformed into the canonical form that is invariant under $O(3)$ space group transformations. With this particular module, the equivariance can be captured, meaning that the subsequent components of GoeCTP are no longer required to account for equivariance.
Crystal Graph Construction. To enable networks to handle such infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is typically necessary to employ graph construction methods that represent the interactions between infinite crystal structures and atoms using finite graph data. Here, we use the crystal graph construction from Yan et al. [50, 51] to describe the structure and relationships within crystals. A detailed explanation of the crystal graph construction can be found in Appendix A.2. Upon construction, the edge features are denoted as $e_{ij}$, while the node features, representing atomic properties, are denoted as $\bm{a}_{i}$.
Node and Edge Feature Embedding. Building on previous work [47, 51, 50], node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$. Edge features $\mathbf{e}_{ij}$ are decomposed into their magnitude $||\mathbf{e}_{ij}||_{2}$ and a normalized direction vector $\hat{\mathbf{e}}_{ij}$. The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding [21]. Subsequently, $\mathbf{e}_{ij}$ are embedded into feature vector $\bm{f}_{ij}^{e}$.
Prediction Module. Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by Yan et al. [50], Taniai et al. [42], Yan et al. [49], Lee et al. [20], Wang et al. [44], among others. For better performance, we select eComformer [50], which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module. A detailed explanation of eComFormer can be found in Appendix A.2. Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows: $\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}$.
Reverse R&R Module. The primary function of the Reverse R&R module is to generate the equivariant tensor property predictions based on the crystal global features from the Prediction module. First, the Reverse R&R module transforms $\bm{G}^{\mathrm{final}}$ into a tensor output, as follows:
$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}})$ (5)
$\bm{\varepsilon}^{\mathrm{final}}=f_{rp}(\bm{\varepsilon},\mathbf{Q}),$ (6)
where $f_{MLP}(\cdot)$ represents a multilayer perceptron (MLP) and the operation reshaping dimension. Next, $f_{rp}(\cdot)$ utilizes the orthogonal matrix $\mathbf{Q}$ obtained from the R&R Module to convert the tensor output $\bm{\varepsilon}$ into its final equivariant form denoted as $\bm{\varepsilon}^{\mathrm{final}}$.
This conversion $f_{rp}(\cdot)$ for predicting the dielectric tensor can be expressed by:
$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}.$ (7)
For predicting the higher-order piezoelectric and elastic tensor, the conversion process becomes more complex, see Appendix A.8 for more details.
4 Experiments
4.1 Experimental setup
Datasets. In this work, we evaluate the performance of GoeCTP on three key tensor property prediction tasks: the second-order dielectric tensor, the third-order piezoelectric tensor, and the fourth-order elastic tensor, respectively. Further details of the dataset and experimental setup can be found in Appendix A.10.
Baseline Methods. We selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e. MEGNET [1, 31], EGTNN [53], and GMTNet [51], as baseline methods.
[…]
4.2 Experimental Results
Predicting Dielectric Tensors. The performance of various models in predicting the dielectric tensor is summarized in Table 1. While GoeCTP and GMTNet show close results in terms of Fnorm and EwT 25%, GoeCTP demonstrates higher values for EwT 10% and EwT 5%, with an improvement of 5% in EwT 5% compared to GMTNet. This indicates that GoeCTP delivers higher-quality predictions compared to GMTNet. Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer [50] and CrystalFormer [42]. Detailed results of these combinations are provided in Appendix A.11. In addition to these experimental results, we further investigate the impact of different canonical forms on tensor property prediction in crystalline materials. Detailed analyses can be found in Appendix A.16.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5%
Table 1: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the dielectric dataset.
Predicting Piezoelectric Tensors. The experimental results for the elastic tensor dataset are shown in Table 2. Similar to the experimental results on the dielectric tensors dataset, it achieves optimal results on all evaluation metrics. Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 0.504 0.515 0.450 0.431
EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9%
EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9%
EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7%
Table 2: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the piezoelectric dataset.
Predicting Elastic Tensors. Experimental results on the elastic tensor dataset are presented in Table 3. GoeCTP demonstrated outstanding performance in predicting higher-order, complex tensors, outperforming all baseline methods across every evaluation metric. Notably, it achieved the lowest Fnorm of 107.11 and improved all EwT metrics by an average of 6% compared to GMTNet, emphasizing its general applicability in predicting diverse tensor properties in materials science. Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2%
Table 3: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the elastic dataset.
eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
Dielectric: Fnorm $\downarrow$ 3.23 4.71 3.23 3.23
Total Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
Piezoelectric: Fnorm $\downarrow$ 0.431 0.527 0.431 0.431
Total Time (s) $\downarrow$ 37.83 37.23 38.01 37.71
Elastic: Fnorm $\downarrow$ 107.11 138.45 107.11 107.11
Total Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric, piezoelectric, and elastic dataset. After training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets
Verifying the $O(3)$ Equivariance. To evaluate the effectiveness of GoeCTP, we conducted experiments to verify the $O(3)$ equivariance of tensor properties. Specifically, after training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set). All crystals in the original test set were adjusted to the canonical form, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set. The method for generating the corresponding orthogonal matrices is from Heiberger [10]. We then evaluated both the Prediction module (eComFormer) and GoeCTP on these two datasets and compared the performance metrics.
The results on the three tensor datasets are shown in Table 4, GoeCTP performed equally well on the augmented test set as on the original test set in all tensor datasets, indicating that it maintains strong $O(3)$ equivariance for tensor properties. In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45% in dielectric tensor dataset, nearly 22% in piezoelectric tensor dataset, and nearly 29% in elastic tensor dataset, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 663 1325 1611 616
Time/batch (s) $\downarrow$ 0.043 0.085 0.103 0.039
Table 5: Efficiency comparison on the dielectric dataset.
Additionally, runtime comparisons between GoeCTP and eComFormer on the test set revealed that GoeCTP did not result in a significant increase in runtime. This demonstrates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality while maintaining efficiency.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 943 1220 5771 938
Time/batch (s) $\downarrow$ 0.060 0.078 0.370 0.060
Table 6: Efficiency comparison on the piezoelectric dataset.
Efficiency. The results presented in Table 5, Table 6, and Table 7 illustrate the running times for GoeCTP and baseline methods. On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet. On the piezoelectric dataset, GoeCTP completed the entire training process with only 16.3% of the time spent compared to GMTNet. On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422
Time/batch (s) $\downarrow$ 0.036 0.056 $>$ 0.452 0.030
Table 7: Efficiency comparison on the elastic dataset.
To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply. In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a MLP at the network’s output, ensuring both efficiency and scalability across different tensor orders.
5 Concluding Remarks
In this work, we revisit crystal tensor property prediction through the lens of canonicalization. Built upon the notion of canonicalization, we proposed a novel $O(3)$-equivariant framework GoeCTP for fast and accurate crystal tensor prediction. Benefiting from the canonicalization mechanism, GoeCTP serves as a plug-and-play module that can be integrated with any existing scalar property prediction network, enabling it to predict tensor properties with negligible additional computational cost. The proposed approach achieves state-of-the-art predictive performance and superior efficiency across a variety of widely used benchmark datasets for crystal tensor properties.
Except for the excel performance, our current GoeCTP has some limitations remain, which include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module. (2) GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types (See Appendix A.12 for more limitations details). In future study, we intend to improve performance by integrating prior knowledge related to the independent components of tensor properties across various crystal systems (see Appendix A.9). Further discussion on the utilization of space group constraints on tensor properties can be found in Appendix A.13.
Appendix A Appendix
A.1 Related work
GNN-Based Methods. CGCNN is a pioneering GNN model specifically designed for handling crystal structures [47]. This model proposed to represent crystal structures as multi-edge crystal graphs. It was applied to predict various single-value properties such as formation energy and band gap. Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge [1, 26, 2, 3, 21]. These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors. Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction. In contrast, recent studies attempted to ensure equivariance through specialized network architectures [28, 25, 11, 46, 34, 51, 53]. These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties. Within these network architectures, many operations are required, such as tensor products and combining irreducible representations. These processes significantly increase computational costs, especially when handling higher-order data.
Transformer-Based Methods. Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties. Matformer [49], one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells. This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures. Subsequently, more advanced Transformer-based approaches were proposed. Some methods, such as ComFormer [50], CrystalFormer [44], and CrystalFormer [42], typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties. Furthermore, DOSTransformer [20] is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance. Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
Equivariant Learning. Recently, two general-purpose equivariant learning techniques for the arbitrary group have gained widespread attention: frame averaging [22, 36, 5] and canonicalization [27, 17, 6]. Frame averaging is developed based on group averaging and involves averaging over smaller frames instead of the entire group. Canonicalization, on the other hand, is inherently related to frame averaging and can be viewed as a more general form of it. [27]. In particular, Lin et al. [22] introduced minimal frame averaging, a method designed to achieve both efficient and exact equivalence in equivariant learning. This approach is primarily applicable to scenarios where the input and output spaces are identical, such as the n-body problem. It does not extend to cases involving different spaces, such as crystal tensor prediction. If this method were extended to tensor property prediction, it would represent a special case of Proposition 3.4, where both orbit canonicalization and rigid registration are achieved using QR decomposition. Within Proposition 3.4, however, different functions can be employed for orbit canonicalization and rigid registration. For instance, the Kabsch algorithm [18, 19] can be employed for rigid registration, while polar decomposition can be utilized for canonicalization.
To the best of our knowledge, no previous work has extended these equivariant learning techniques to the prediction of crystal tensor properties. Thus, our method represents the first extension of these concepts to the crystal tensor property prediction.
A.2 Details of Framework Architecture
Crystal Graph Construction. In GoeCTP, we use the crystal graph construction from Yan et al. [50, 51] to describe the structure and relationships within crystals. Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{X},\mathbf{H})$ as introduced in Section 2. Then, we assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$. $\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$. When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e. $d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$. Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$. Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al. [44], Yan et al. [50].
eComFormer. The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks [50]. It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph. The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. [50].
[…]
$\bm{\alpha}_{ij}=\frac{\bm{q}_{ij}\circ\bm{\xi}_{K}(\bm{k}_{ij})}{\sqrt{d_{\bm{q}_{ij}}}},\bm{msg}_{ij}=\operatorname{sigmoid}(\operatorname{BN}(\bm{\alpha}_{ij}))\circ\bm{\xi}_{V}(\bm{\upsilon}_{ij}),$ (8)
where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively. The operators $\circ$ denote the Hadamard product, $\operatorname{BN}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$. Then, node feature $\bm{f}_{i}$ is updated as follows,
$\bm{m}s\bm{g}_{i}=\sum_{j\in\mathcal{N}_{i}}\bm{m}s\bm{g}_{ij},\bm{f}_{i}^{\mathrm{new}}=\bm{\xi}_{msg}(\bm{f}_{i}+\mathrm{BN}(\bm{m}s\bm{g}_{i})),$ (9)
where $\bm{\xi}_{msg}(\cdot)$ denoting the softplus activation function.
The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers [7]. It uses node feature $\bm{f}_{i}^{l}$ and equivalent vector feature $\mathbf{e}_{ji}$ embedded by corresponding spherical harmonics $\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})=c_{0}$,$\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})=c_{1}*\frac{\mathbf{e}_{ji}}{||\mathbf{e}_{ji}||_{2}}\in\mathbb{R}^{3}$ and $\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{5}$ to represent the input features. Gathering rotational information from neighboring nodes to the central node $i$, the first TP layer is shown as
$\bm{f}_{i,0}^{l}=\bm{f}_{i}^{l^{\prime}}+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})),\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),\lambda\in\{1,2\},$ (10)
where $\bm{f}_{i}^{l^{\prime}}$ is linearly transformed from $\bm{f}_{i}^{l}$, $|\mathcal{N}_{i}|$ and $\mathbf{TP}_{\lambda}$ represent the number of neighbors of node $i$ and TP layers with rotation order $\lambda$ respectively. Then, to represent the invariant node features, the second TP layer is written as
$\bm{f}_{i}^{l*}=\frac{1}{|\mathcal{N}_{i}|}(\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,0}^{l},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,1}^{l},\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,2}^{l},\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji}))),$ (11)
stacking the two tensor product layers together using both linear and nonlinear transformations, the output $\bm{f}_{i,updated}^{l}$ is combined as
$\bm{f}_{i,updated}^{l}=\bm{\sigma}_{\mathrm{equi}}(\mathrm{BN}(\bm{f}_{i}^{l*}))+\mathrm{LN}_{\mathrm{equi}}(\bm{f}_{i}^{l}),$ (12)
with $\bm{\sigma}_{\mathrm{equi}}$ denoting a nonlinear transformation made up of two softplus layers with a linear layer positioned between them. The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer are shown in Fig. 3.
A.3 Proofs
A.4 Proofs of Lemma A.1
Lemma A.1.
Given a orbit canonicalization function $C_{\mathbf{M}}(\cdot)$, for any prediction function ${f_{\theta}}(\cdot)$, it holds that ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ is $O(3)$-invariant.
Proof.
According to Definition 3.2, for any $g\in O(3)$ and $\mathbf{M}\in\mathcal{V}$, we have $C_{\mathbf{M}}(\mathbf{M})=\mathbf{M_{0}}$ and $C_{\mathbf{M}}(g\cdot\mathbf{M})=\mathbf{M_{0}}$. Thus,
${f_{\theta}}(C_{\mathbf{M}}(g\cdot\mathbf{M}))$ $={f_{\theta}}(\mathbf{M_{0}})$ (13)
$={f_{\theta}}(C_{\mathbf{M}}(\mathbf{M})),$
which implies ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ is $O(3)$-invariant. ∎
A.5 Proofs of Lemma A.2
Lemma A.2.
For a rigid registration function $R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})=g$, it is $O(3)$-equivariant with respect to the input $\mathbf{M}_{1}$.
Proof.
According to Definition 3.3, $\mathbf{M}_{1}=R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$. Thus, for any $g\in O(3)$, we have
$R_{\mathbf{M}}(g\cdot\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$ $=g\cdot\mathbf{M}_{1}$ (14)
$=g\cdot(R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2})$
$=(g\cdot R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2}))\cdot\mathbf{M}_{2}.$
Given $\mathbf{M}_{2}=(\mathbf{A}_{2},\mathbf{F}_{2},\mathbf{L}_{2})$, and noting that $g\cdot\mathbf{M}_{2}=(\mathbf{A}_{2},\mathbf{F}_{2},\mathbf{Q}\mathbf{L}_{2})$ (where $\mathbf{L}_{2}$ has full rank), we can write Eq. 14 as
$(R_{\mathbf{M}}(g\cdot\mathbf{M}_{1},\mathbf{M}_{2}))\mathbf{L}_{2}=(g\cdot R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2}))\mathbf{L}_{2}.$ (15)
Therefore, $R_{\mathbf{M}}(g\cdot\mathbf{M}_{1},\mathbf{M}_{2})=g\cdot R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})$, which implies that $R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})$ is $O(3)$-equivariant. ∎
A.6 Proof of Proposition 3.4
Proposition 3.4. ($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$ equivariant for tensor prediction.
Proof.
According to Lemma A.2 and Lemma A.1, for any $g\in O(3)$ we have
$h(g\cdot\mathbf{M})$ $=R_{\mathbf{M}}(g\cdot\mathbf{M},C_{\mathbf{M}}(g\cdot\mathbf{M}))\cdot f(C_{\mathbf{M}}(g\cdot\mathbf{M}))$ (16)
$=R_{\mathbf{M}}(g\cdot\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$
$=g\cdot R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$
$=g\cdot h(\mathbf{M})$
Thus, $h(\mathbf{M})$ is $O(3)$ equivariant for tensor prediction.
A.7 Proof of Proposition 3.6
Proposition 3.6. (Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix. In this perspective, $f_{p1}(\cdot)$ serve as the orbit canonicalization function, while $f_{p2}(\cdot)$ serve as the rigid registration function. The matrix $\mathbf{H}$ is thus identified as the canonical form.
Proof.
Given the lattice matrices $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$, where $\mathbf{Q^{\prime}}$ is an arbitrary $O(3)$ group transformation, applying polar decomposition on both $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$ yields the following expressions:
$\mathbf{L}=\mathbf{Q}\mathbf{H}$ (17)
$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (18)
By combining Eq. 17 and 18, we can have the following equation:
$\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (19)
Because $\mathbf{Q^{\prime}}\mathbf{Q}(\mathbf{Q^{\prime}}\mathbf{Q})^{\top}=\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{Q}^{\top}\mathbf{Q^{\prime}}^{\top}=\mathbf{I}$, it follows that $\mathbf{Q^{\prime}}\mathbf{Q}$ is an orthogonal matrix. According to the uniqueness of polar decomposition, we can have $\mathbf{Q^{\prime}}\mathbf{Q}=\mathbf{Q_{1}}$, which implies $\mathbf{Q_{1}}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}$. If $\mathbf{H_{1}}\neq\mathbf{H}$, this would violate the uniqueness property of polar decomposition. Therefore, we conclude that $\mathbf{H_{1}}=\mathbf{H}$, indicating that $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation. Consequently, the crystal configuration $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$ can serve as the canonical form. Performing polar decomposition on the lattice matrix to obtain $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$, i.e. $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$, can thus be regarded as the orbit canonicalization function. Furthermore, based on Eq. 17, performing polar decomposition on the lattice matrix to obtain $\mathbf{Q}$, i.e. $f_{p2}(\mathbf{M},\mathbf{M_{0}})=\mathbf{Q}$, can be regarded as rigid registration function, as shown below:
$\mathbf{L}=\mathbf{Q}\mathbf{H}\quad\to\quad\mathbf{M}=\mathbf{Q}\mathbf{M_{0}}\quad\to\quad\mathbf{M}=f_{p2}(\mathbf{M},\mathbf{M_{0}})\cdot\mathbf{M_{0}}.$ (20)
A.8 $O(3)$ equivariance for crystal tensor properties
Due to the difference between the input space $\mathcal{V}$ and the output space $\mathcal{W}$, the $O(3)$ equivariance required for crystal tensor property prediction tasks differs from the $O(3)$ equivariance typically encountered in general molecular studies [13, 48, 40]. In molecular studies, the input space $\mathcal{V}$ and the output space $\mathcal{W}$ are the same, meaning group representations in these spaces are identical, i.e. $\rho_{\mathcal{V}}(g)=\rho_{\mathcal{W}}(g)$. Specifically, for a function $f(\mathbf{X})$, if it is $O(3)$ equivariant, then it satisfies $f(\mathbf{Q}\mathbf{X})=\mathbf{Q}f(\mathbf{X})$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an arbitrary orthogonal matrix, and $\mathbf{X}\in\mathbb{R}^{3\times N}$ represents the coordinate matrix of atoms in a molecule. In this context, the function $f$ can be seen as the model. However, in the prediction of crystal tensor properties, the requirements for $O(3)$ equivariance differ. In what follows, we will provide a detailed introduction to this difference.
Dielectric tensor. For the dielectric tensor $\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$ and a function $f:(\mathbf{A},\mathbf{F},\mathbf{L})\to\bm{\varepsilon}$, if it is $O(3)$ equivariant, it must satisfy $f(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top}$. The reason for this difference lies in the physical nature of the dielectric tensor, which characterizes a material’s polarization response to an external electric field [51]. Specifically, the dielectric tensor describes the relationship between the electric displacement $\mathbf{D}\in\mathbb{R}^{3}$ and the applied electric field $\mathbf{E}\in\mathbb{R}^{3}$ through the equation $\mathbf{D}=\bm{\varepsilon}\mathbf{E}$. When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal structure, the relationship $\mathbf{D^{\prime}}=\bm{\varepsilon}^{\prime}\mathbf{E^{\prime}}$ holds, where $\mathbf{D^{\prime}}=\mathbf{Q}\mathbf{D}$ and $\mathbf{E^{\prime}}=\mathbf{Q}\mathbf{E}$. Substituting these into the equation, we have:
$\mathbf{Q}\mathbf{D}=\bm{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E}\to\mathbf{D}=\mathbf{Q}^{\top}\bm{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E},$ (21)
which implies that the dielectric tensor transforms under the $O(3)$ group transformation as: $\bm{\varepsilon}^{\prime}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}$. This transformation principle extends similarly to other crystal tensors, as follows.
Piezoelectric tensor. The piezoelectric tensor $\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$ describes the relationship between the applied strain $\bm{\epsilon}\in\mathbb{R}^{3\times 3}$ to the electric displacement field $\mathbf{D}\in\mathbb{R}^{3}$ within the material. Mathematically, this relationship is expressed as $\mathbf{D}_{i}=\sum_{jk}\mathbf{e}_{ijk}\bm{\epsilon}_{jk}$, with $i,j,k\in\{1,2,3\}$.
When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and electric displacement field is transformed to $\bm{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}_{mn}$ and $\mathbf{D}_{i}^{\prime}=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$. The relation is then reformulated as $\mathbf{D}_{i}^{\prime}=\sum_{jk}\mathbf{e}_{ijk}^{\prime}\bm{\epsilon}_{jk}^{\prime}$. Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\bm{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn}$. Consequently, $\mathbf{D}_{i}^{\prime}$ can be represented as
$\mathbf{D}_{i}^{\prime}$ $=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$ (22)
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}\bm{\epsilon}_{jk}$
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}(\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn})$
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{mn}\mathbf{e}_{\ell mn}(\sum_{jk}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}^{\prime}_{jk})\quad(exchange\,sign,m\leftrightarrow j,n\leftrightarrow k)$
$=\sum_{jk}\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}\bm{\epsilon}_{jk}^{\prime}$
Therefore, under the $O(3)$ group transformation $\mathbf{Q}$, the transformed piezoelectric tensor $\mathbf{e}_{ijk}^{\prime}$ is given by:
$\mathbf{e}_{ijk}^{\prime}=\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}.$ (23)
When employing GoeCTP to predict the piezoelectric tensor, Equation 7 becomes $\mathbf{e}_{ijk}^{\mathrm{final}}=\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}$.
Elastic tensor. The elastic tensor $C\in\mathbb{R}^{3\times 3\times 3\times 3}$ describes the relationship between the applied strain $\bm{\epsilon}\in\mathbb{R}^{3\times 3}$ and the stress tensor $\sigma\in\mathbb{R}^{3\times 3}$ within the material. This relationship is expressed as $\bm{\sigma}_{ij}=\sum_{k\ell}C_{ijk\ell}\bm{\epsilon}_{k\ell}$, with $i,j,k,\ell\in\{1,2,3,4\}$.
When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and stress tensor is transformed to $\bm{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}_{mn}$ and $\bm{\sigma}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\sigma}_{mn}$. Under this transformation, the relationship becomes $\bm{\sigma}_{ij}^{\prime}=\sum_{k\ell}C_{ijk\ell}^{\prime}\bm{\epsilon}_{k\ell}^{\prime}$. Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\bm{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn}$. Based on above equations, the transformed stress tensor $\bm{\sigma}_{ij}^{\prime}$ can be expressed as:
$\bm{\sigma}_{ij}^{\prime}$ $=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\bm{\sigma}_{mn}$ (24)
$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\bm{\epsilon}_{pq}$
$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\sum_{k\ell}\mathbf{Q}_{kp}\mathbf{Q}_{\ell q}\bm{\epsilon}^{\prime}_{k\ell}$
$=\sum_{k\ell}\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}\bm{\epsilon}^{\prime}_{k\ell}$
Therefore, under the $O(3)$ group transformation $\mathbf{Q}$, $C_{ijkl}^{\prime}$ is represented as:
$C_{ijkl}^{\prime}=\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}.$ (25)
When GoeCTP is used to predict elastic tensor, Equation 7 becomes $C_{ijkl}^{\mathrm{final}}=\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}$.
A.9 Tensor properties symmetry
Tensor properties describe the material’s response to external physical fields (such as electric fields or mechanical stress). In materials with symmetry, this response must adhere to the symmetry requirements of the material. As demonstrated by Yan et al. [51], when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$. Therefore, the corresponding tensor properties also remain unchanged, i.e., $\bm{\varepsilon}=\mathbf{R}\bm{\varepsilon}\mathbf{R}^{\top}$. Thus, crystal symmetry imposes strict constraints on the components of the tensor, leading to the simplification or elimination of many components, reducing the number of independent components.
Crystal system Number of independent elements Dielectric tensor
Cubic 1 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Tetragonal & Hexagonal & Trigonal 2 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Orthorhombic 3 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{22}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Monoclinic 4 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&\bm{\varepsilon}_{13}\\ 0&\bm{\varepsilon}_{22}&0\\ \bm{\varepsilon}_{13}&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Triclinic 6 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&\bm{\varepsilon}_{12}&\bm{\varepsilon}_{13}\\ \bm{\varepsilon}_{12}&\bm{\varepsilon}_{22}&\bm{\varepsilon}_{23}\\ \bm{\varepsilon}_{13}&\bm{\varepsilon}_{23}&\bm{\varepsilon}_{33}\end{pmatrix}$
Table 8: Number of independent components in the dielectric tensor for different crystal systems.
Independent components in the dielectric tensor. The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry. The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 8 [28].
Voigt notation for elastic tensor. Similar to the dielectric tensor, the elastic tensor also has independent elements for various types of systems due to the crystal symmetry. The elastic tensor has a minimum of 3 and a maximum of 21 independent elements for various types of systems. Voigt notation is a compact way to represent these independent components of tensor properties [14]. According to the rules $11\to 1\mathrm{~{};~{}}22\to 2\mathrm{~{};~{}}33\to 3\mathrm{~{};~{}}23,~{}32\to 4\mathrm{~{};~{}}31,~{}13\to 5\mathrm{~{};~{}}12,~{}21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix [46, 37]:
$C=\begin{pmatrix}C_{1111}&C_{1122}&C_{1133}&C_{1123}&C_{1131}&C_{1112}\\ C_{1122}&C_{2222}&C_{2233}&C_{2223}&C_{2231}&C_{2212}\\ C_{1133}&C_{2233}&C_{3333}&C_{3323}&C_{3331}&C_{3312}\\ C_{1123}&C_{2223}&C_{3323}&C_{2323}&C_{2331}&C_{2312}\\ C_{1131}&C_{2231}&C_{3331}&C_{2331}&C_{3131}&C_{3112}\\ C_{1112}&C_{2212}&C_{3312}&C_{2312}&C_{3112}&C_{1212}\end{pmatrix}\to\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}$
The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to Wen et al. [46], Ran et al. [37]).
Crystal system Number of independent elements Elastic tensor
Cubic 3 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{12}&0&0&0\\ C_{12}&C_{11}&C_{12}&0&0&0\\ C_{12}&C_{12}&C_{11}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{44}&0\\ 0&0&0&0&0&C_{44}\end{pmatrix}\end{gathered}$
Tetragonal 6 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&0&0&0\\ C_{12}&C_{11}&C_{13}&0&0&0\\ C_{13}&C_{13}&C_{33}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{44}&0\\ 0&0&0&0&0&C_{66}\end{pmatrix}\end{gathered}$
Triclinic 21 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}\end{gathered}$
Table 9: Number of independent components in the elastic tensor for different crystal systems.
Voigt notation for piezoelectric tensor. The number of independent components in the piezoelectric tensor for different crystal systems is shown in Tabel 10 (partial data shown; for more details, refer to De Jong et al. [4], Gorfman and Zhang [8]).
Crystal system point groups Number of independent elements Piezoelectric tensor
Trigonal 32 2 $\mathbf{e}=\begin{pmatrix}\mathbf{e}_{11}&-\mathbf{e}_{11}&0&\mathbf{e}_{14}&0&0\\ 0&0&0&0&-\mathbf{e}_{14}&-\mathbf{e}_{11}\\ 0&0&0&0&0&0\end{pmatrix}$
Monoclinic 2 8 $\mathbf{e}=\begin{pmatrix}0&0&0&\mathbf{e}_{14}&0&\mathbf{e}_{16}\\ \mathbf{e}_{21}&\mathbf{e}_{22}&\mathbf{e}_{23}&0&e_{25}&0\\ 0&0&0&\mathbf{e}_{34}&0&\mathbf{e}_{36}\end{pmatrix}$
Triclinic 1 18 $\mathbf{e}=\begin{pmatrix}\mathbf{e}_{11}&\mathbf{e}_{12}&\mathbf{e}_{13}&\mathbf{e}_{14}&\mathbf{e}_{15}&\mathbf{e}_{16}\\ \mathbf{e}_{21}&\mathbf{e}_{22}&\mathbf{e}_{23}&\mathbf{e}_{24}&\mathbf{e}_{25}&\mathbf{e}_{26}\\ \mathbf{e}_{31}&\mathbf{e}_{32}&\mathbf{e}_{33}&\mathbf{e}_{34}&\mathbf{e}_{35}&\mathbf{e}_{36}\end{pmatrix}$
Table 10: Number of independent components in the piezoelectric tensor for different crystal systems.
A.10 Experimental details
Datasets details. The dataset for dielectric tensor and piezoelectric tensor is derived from the data processed by Yan et al. [51], sourced from the JARVIS-DFT database. Similarly, the dataset for elastic tensor is obtained from the dft_3d data within the JARVIS-DFT database111 https://pages.nist.gov/jarvis/databases/. The statistical details of the datasets are presented in Table 11. Additionally, in the dielectric tensor dataset, the dielectric tensor is a $3\times 3$ symmetric matrix. Therefore, during prediction, we predict 6 elements of the matrix (see Appendix A.9 for details) and then reconstruct the entire $3\times 3$ symmetric matrix. In the piezoelectric tensor dataset, the piezoelectric tensor is represented using Voigt notation as a $3\times 6$ matrix, rather than a $3\times 3\times 3$ third-order tensor (see Appendix A.9 for details). we only predict the $3\times 6$ matrix. In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor [14, 46]. Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.9 for details).
Dataset Sample size Fnorm Mean Fnorm STD Unit
Dielectric 4713 14.7 18.2 Unitless
Piezoelectric 4998 0.43 3.09 $\text{C}/\text{m}^{2}$
Elastic 25110 306.4 238.4 GPa
Table 11: Dataset statistics.
Experimental Settings. The experiments were performed using a single NVIDIA GeForce RTX 3090 GPU. For benchmarking, we directly utilized the codebases for MEGNET, ETGNN, and GMTNet as provided by Yan et al. [51]. For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio. During model training, Huber loss, AdamW [24], a weight decay of $10^{-5}$, and polynomial learning rate decay were employed.
Hyperparameter settings of GoeCTP. When constructing the crystal graph, we used the 16th nearest atom to determine the cutoff radius. For edge embeddings, we used an RBF kernel with $c=0.75$ and values ranging from --4 to 0, which maps $-c/||\mathbf{e}_{ij}||_{2}$ to a 512-dimensional vector. In the dielectric and piezoelectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector. For the prediction module, eComFormer, in the dielectric and piezoelectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer. For both dielectric, piezoelectric, and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64. In the Reverse R&R module, for the dielectric tensor prediction task, the 128-dimensional vector features output by the prediction module are mapped to a 6-dimensional vector through two non-linear layers, then reconstructed into a $3\times 3$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ obtained from the R&R module to achieve $O(3)$ equivariant output. For the piezoelectric prediction task, the 18-dimensional vector features output by the prediction module are reconstructed into a $3\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output. For the elastic tensor prediction task, the 256-dimensional vector features output by the prediction module are mapped to a 36-dimensional vector through two non-linear layers, then reconstructed into a $6\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
Hyperparameter settings of GMTNet, ETGNN, and MEGNET. Following Yan et al. [51], we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks. The same polynomial learning rate decay scheduler is used in all experiments.
A.11 Additional results.
We assessed the performance of GoeCTP when combined with iComFormer [50] and CrystalFormer [42] (denoted as GoeCTP (iCom.) for the combination) on both the dielectric, piezoelectric, and elastic tensor datasets. As presented in Table 12, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). As shown in Table 13, similar to the dielectric tensor dataset, GoeCTP (eCom.) achieved the best prediction accuracy on the piezoelectric dataset, but GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). In contrast, as shown in Table 14, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23 3.40 3.53
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2% 81.7% 80.1%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8% 53.8% 52.9%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5% 32.3% 30.6%
Total Time (s) $\downarrow$ 663 1325 1611 616 535 645
Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048 0.042 0.202
Table 12: Additional comparison of performance metrics on dielectric dataset.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 0.504 0.515 0.450 0.431 0.448 0.457
EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9% 45.5% 42.7%
EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9% 44.9% 42.1%
EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7% 38.3% 41.5%
Total Time (s) $\downarrow$ 943 1220 5771 938 909 2787
Time/batch (s) $\downarrow$ 0.074 0.095 0.45 0.073 0.071 0.218
Table 13: Additional comparison of performance metrics on piezoelectric dataset.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11 102.80 107.44
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5% 46.7% 43.5%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3% 18.6% 15.8%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2% 8.2% 7.9%
Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422 4035 7891
Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189 0.315 0.616
Table 14: Additional comparison of performance metrics on elastic dataset.
A.12 The limitations of our method and its broader impact
Limitations. (1)As discussed in the main text, GoeCTP is a plug-and-play $O(3)$-equivariant framework designed to enhance the backbone network’s ability to achieve equivariant predictions. Consequently, the performance of our method is inherently dependent on the capabilities of the backbone network. If the backbone lacks sufficient predictive power, the combined framework may not fully surpass the state-of-the-art tensor prediction networks in all evaluation metrics. For instance, as shown in Table 12, when our framework is integrated with CrystalFormer [42], although the computational speed significantly exceeds that of GMTNet, the Fnorm, EwT 25%, and EwT 10% values are lower than those achieved by GMTNet.
(2)As demonstrated in Higham [12] and Proposition 3.5, polar decomposition applied to a $3\times 3$ invertible matrix $\mathbf{L}$ produces a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$. For 3D crystal structures, the lattice matrix $\mathbf{L}$ is always full-rank (i.e. invertible), therefore, ensuring the applicability of our method to 3D crystal systems. However, for certain special cases, such as 2D crystals with single-layer structures [32, 39], the rank of the lattice matrix $\mathbf{L}$ may be less than 3. In these scenarios, directly applying polar decomposition may not yield a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$ and a unique canonical form $\mathbf{H}$. This limitation could cause our method to fail, as crystals with different spatial orientations cannot be consistently adjusted to a unique canonical form, thereby preventing the method from achieving the desired $O(3)$-equivariance.
Broader impact. This work can be applied to screen existing crystal material databases for potential material candidates, facilitating the discovery of new materials with desirable tensor properties, such as high-dielectric materials. Therefore, our work has the potential to make a significant impact on materials science.
A.13 How to further utilize the tensor properties symmetry
In this section, we will use the dielectric tensor as an example to briefly discuss how to further utilize the symmetry of tensor properties.
Utilizing symmetry constraints for zero elements. To demonstrate the role of symmetry constraints, we present an example of the GoeCTP prediction results in Table 15. Within a reasonable margin of error, the predictions align well with the expected constraints.
Label Prediction Cubic dielectric tensor
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.252&0.016&0.008\\ 0.016&2.230&0.007\\ 0.008&0.007&2.262\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Table 15: An example of GoeCTP prediction results
Crystal system Cubic Tetragonal Hexagonal-Trigonal Orthorhombic Monoclinic
Success rate 88.3% 86.6% 84.5% 84.5% 75.7%
Table 16: The GoeCTP results of predicting symmetry-constrained zero-valued dielectric tensor elements.
For a dielectric tensor, we evaluate the prediction of zero elements by defining a threshold as 1% of the average value of non-zero elements in the labels. Predictions are deemed successful if they meet this threshold. The results are as shown in Table 16. It can be observed that our method predicts most zero elements accurately, though not flawlessly. Since our advantage lies in transferring equivariance through an external framework, there are no restrictions on the model itself. Therefore, to achieve a higher success rate for zero elements, we added a ReLU activation function to the output layer of the network to improve the success rate (this applies only to cases where tensor elements are greater than or equal to zero). The results after retraining GoeCTP are as shown in Table 17, Table 18, and Table 19. Our method predicts zero elements more accurately while maintaining overall performance.
Label Prediction Cubic dielectric tensor
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.237&0.000&0.000\\ 0.000&2.283&0.000\\ 0.000&0.000&2.228\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Table 17: An example of GoeCTP (ReLU) prediction results
Crystal system Cubic Tetragonal Hexagonal-Trigonal Orthorhombic Monoclinic
Success rate 100% 100% 87.2% 100% 100%
Table 18: The GoeCTP (ReLU) results of predicting symmetry-constrained zero-valued dielectric tensor elements.
GoeCTP GoeCTP (ReLU)
Fnorm $\downarrow$ 3.23 3.26
EwT 25% $\uparrow$ 83.2% 82.6%
EwT 10% $\uparrow$ 56.8% 58.4%
EwT 5% $\uparrow$ 35.5% 36.3%
Table 19: Performance comparisons between GoeCTP and GoeCTP (ReLU) on the dielectric dataset.
This simple modification enables GoeCTP to more accurately predict the zero elements in dielectric tensors caused by the space group constraints.
Utilizing symmetry constraints for non-zero elements. The symmetry constraints for zero elements serve as a foundational example. When the space group of the input crystal is known, prior knowledge (refer to Appendix A.9) can be employed to ensure that network outputs fully adhere to tensor property constraints. There are two specific methods to achieve this.
In the first method, only the independent components of the tensor properties are predicted. For example, for a cubic crystal, we predict only one independent component and then use the mask from Table 8 to reconstruct it into a $3\times 3$ tensor). In Table 15, GoeCTP was modified to directly predict independent components. The new results are shown in Table 20, which fully satisfy the symmetry constraints. As shown in Table 21, the overall prediction performance indicates that directly predicting the independent components improves the model’s Fnorm performance on the dielectric dataset. Since this method directly predicts independent components, it is more suitable for cases like the dielectric tensor, where the independent components exhibit relatively simple structures. However, for more complex datasets, such as those involving piezoelectric and elastic tensors, the model’s performance using this method may degrade. In such scenarios, the second method provides a more viable alternative.
Label Prediction Cubic dielectric tensor
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.357&0.000&0.000\\ 0.000&2.357&0.000\\ 0.000&0.000&2.357\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Table 20: An example of GoeCTP (predicting the independent components) prediction results
GoeCTP GoeCTP (independent)
Fnorm $\downarrow$ 3.23 3.11
EwT 25% $\uparrow$ 83.2% 80.5%
EwT 10% $\uparrow$ 56.8% 57.3%
EwT 5% $\uparrow$ 35.5% 35.5%
Table 21: Performance comparisons between GoeCTP and GoeCTP (predicting the independent components) on the dielectric dataset.
In the second method, once the network training is completed, the network output is weighted using the standard tensor property format, referred to as a mask, as detailed in Table 8. The application of this method to the piezoelectric and elastic datasets is demonstrated in Tables 22 and 23. The results indicate that incorporating symmetry-based weighting effectively improves the performance of GoeCTP.
GoeCTP GoeCTP (weighted)
Fnorm $\downarrow$ 0.431 0.389
EwT 25% $\uparrow$ 46.9% 48.5%
EwT 10% $\uparrow$ 42.9% 46.7%
EwT 5% $\uparrow$ 39.7% 46.1%
Table 22: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the piezoelectric dataset.
GoeCTP GoeCTP (weighted)
Fnorm $\downarrow$ 107.11 106.24
EwT 25% $\uparrow$ 42.5% 46.3%
EwT 10% $\uparrow$ 15.3% 16.5%
EwT 5% $\uparrow$ 7.2% 7.4%
Table 23: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the elastic dataset.
A.14 Analysis of discontinuity in canonicalization
Previous studies [6] have shown that canonicalization can often introduce discontinuities into the model. To assess the impact of such discontinuities on our method, we analyze our proposed method from both theoretical and empirical perspectives.
From theoretical perspective: The following proposition, presented in Appendix B of Dym et al. [6], characterizes the existence of a continuous canonicalization under orthogonal group actions:
Proposition A.3.
Consider the action of the orthogonal group $O(d)$ on $\mathbf{X}\in\mathbb{R}^{n\times d}$. There exists a continuous canonicalization for $O(d)$ when $n\leq d$. This continuous canonicalization can be defined as $y_{\mathbf{X}}=\left(\mathbf{X}^{\top}\mathbf{X}\right)^{1/2}$, where $\mathbf{X}^{\top}\mathbf{X}$ forms a positive semi-definite matrix and the square root is the standard square root of a positive semi-definite matrix, i.e. $\left(\mathbf{A}\right)^{1/2}$ is the unique positive semidefinite matrix $\mathbf{B}$ such that $\mathbf{B}^{2}=\mathbf{A}$.
In fact, the polar decomposition adopted in this work can be regarded as a case of such continuous canonicalization.
Proof.
According to Proposition A.3, for a matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$, its continuous canonicalization norm is given by $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}$.
In this work, the polar decomposition is applied to the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H},\mathbf{L}\in\mathbb{R}^{3\times 3}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{H}$ is the positive semi-definite Gram matrix (which serves as the canonicalization norm). Based on the properties of positive semi-definite matrices, our method is actually equivalent to the canonicalization approach described in Proposition A.3, i.e. $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}=\left((\mathbf{Q}\mathbf{H})^{\top}\mathbf{Q}\mathbf{H}\right)^{1/2}=\left(\mathbf{H}^{\top}\mathbf{H}\right)^{1/2}=\mathbf{H}$. ∎
Therefore, our canonicalization via polar decomposition is consistent with the continuous canonicalization.
From empirical perspective: To further demonstrate how this discontinuity affects the accuracy of tensor property predictions, we conducted a set of comparative experiments in which we replaced the polar decomposition in GoeCTP with QR decomposition [22] for evaluation. In this QR decomposition case, the lattice matrix is decomposed as $\mathbf{L}=\mathbf{Q}\mathbf{R}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{R}$ is an upper triangular matrix that serves as the canonicalization norm. However, $\mathbf{R}$ does not satisfy Proposition A.3, and therefore does not constitute a continuous canonicalization approach.
To assess the sensitivity of each approach to small structural perturbations, we used the crystal $CdI_{2}$ (JARVIS_ID: JVASP-29631) as the model input, and then perturbed its lattice matrix (specifically, scaling down the first row of the lattice matrix) to observe the relative changes in the output (specifically the variations in the Fabenius norm of the output tensor). The detailed results are as follows:
Perturbation Ratio 0% 5% 10% 15% 20% 25% 30% 35% 40%
Output variation ratio (%): GoeCTP (QR) 0 3.38 6.63 11.29 21.83 39.37 67.81 124.79 191.28
Output variation ratio (%): GoeCTP (polar) 0 2.53 1.23 2.03 4.84 12.72 24.88 36.34 45.67
Table 24: Output variation ratio under different perturbation ratios on the dielectric dataset
The results clearly show that the output of GoeCTP using polar decomposition exhibits significantly smaller relative changes compared to the QR-based variant, particularly under increasing levels of perturbation. This highlights the robustness of the continuous canonicalization afforded by polar decomposition, demonstrating its superiority in maintaining prediction stability and structural consistency under small input variations.
A.15 How to design other canonical forms
In fact, the canonical form obtained via polar decomposition provides a foundation from which infinitely many alternative canonical forms can be derived. Specifically, given a crystal with lattice matrix $\mathbf{L}$, polar decomposition yields $\mathbf{L}=\mathbf{Q}\mathbf{H}$. While $\mathbf{H}$ serves as one valid canonical representation, additional canonical forms can be constructed by applying orthogonal transformations to $\mathbf{H}$. For instance, given a fixed orthogonal matrix $\mathbf{Q}_{0}$ (which may be arbitrarily chosen) or a neural network $f_{Q}(\mathbf{H})=\mathbf{Q}_{\theta}$ that predicts an orthogonal matrix $\mathbf{Q}_{\theta}$, one can construct new canonical forms such as $\mathbf{Q}_{0}\mathbf{H}$ or $\mathbf{Q}_{\theta}\mathbf{H}$. This is because, for any given crystal with an arbitrary orientation, a fixed canonical form can be obtained according to a predefined computation rule, i.e. $\mathbf{L}=\mathbf{Q}\mathbf{H}\quad\to\quad\mathbf{L}=\mathbf{Q}\mathbf{Q}_{0}^{\top}\mathbf{Q}_{0}\mathbf{H}\quad\text{or}\quad\mathbf{L}=\mathbf{Q}\mathbf{Q}_{\theta}^{\top}\mathbf{Q}_{\theta}\mathbf{H}$. This formulation highlights the inherent flexibility in defining canonical forms.
A.16 Impact of different canonical form on tensor property prediction
As discussed in Appendix A.15, there exist infinitely many possible canonical forms. To investigate the impact of different canonical forms on tensor property prediction in crystalline materials, we first present the results comparing the performance of GoeCTP with QR decomposition [22] and GoeCTP without canonicalization. The corresponding results are summarized in the table below:
GoeCTP (QR) GoeCTP (polar) GoeCTP (w/o QR or polar)
Fnorm $\downarrow$ 3.20 3.23 3.60
EwT 25% $\uparrow$ 83.5% 83.2% 80.6%
EwT 10% $\uparrow$ 56.0% 56.8% 56.2%
EwT 5% $\uparrow$ 32.4% 35.5% 32.6%
Total Time (s) $\downarrow$ 639 616 613
Table 25: Predictive performance comparisons between GoeCTP (QR), GoeCTP (polar), and GoeCTP (w/o QR or polar) on the dielectric dataset.
According to the results, it is evident that incorporating a canonical form improves tensor prediction performance compared to omitting it. While different canonical forms lead to slight variations in performance, the overall impact remains relatively modest. One possible explanation is that canonical forms typically exhibit structured representations, such as the symmetric positive semi-definite matrix obtained via polar decomposition or the upper triangular matrix from QR decomposition, which may facilitate the model’s ability to capture relevant features.
Furthermore, the use of a canonical form aligns all atoms within a crystal structure to a unified global coordinate system, thereby enabling equivariant tensor prediction. However, recent studies [43, 23, 15] have demonstrated that employing adaptive local frames, i.e. constructed based on individual atomic environments, can further enhance the accuracy of property predictions. Motivated by these insights, we extend the GoeCTP framework by incorporating local frames [43] (also referred to as local canonicalization). It is important to emphasize that GoeCTP can be seamlessly integrated with any prediction network, including those based on local frames, while preserving its equivariant prediction capability. The corresponding results are presented below.
GoeCTP (local frame+polar) GoeCTP (polar)
Fnorm $\downarrow$ 3.11 3.23
EwT 25% $\uparrow$ 83.9% 83.2%
EwT 10% $\uparrow$ 58.4% 56.8%
EwT 5% $\uparrow$ 33.5% 35.5%
Total Time (s) $\downarrow$ 1979 616
Table 26: Predictive performance comparisons between GoeCTP (local frame+polar) and GoeCTP (polar) on the dielectric dataset.
The results indicate that incorporating local frames can enhance predictive performance. However, this improvement comes at the expense of increased computational cost.